On Lake Superior there are several islands large enough to support woodland caribou populations. One such island is Michipicoten Island, where caribou populations have been estimated at anywhere from 9 to 100 individuals per year, depending on the strength of wolf predation. Caribou are strong swimmers and can enter and leave the island from the mainland, though they do so rarely. Let’s say there are 90 caribou on the island, and each generation 5 of those 90 caribou emigrated from the mainland. One year the Canadian government performed a genetic census and they determined that the frequency of the Est1 allele (p) was 0.3 on the island, but 0.8 on the neighboring mainland. Using the information provided here and assuming a continent-island model of migration, estimate the frequency of the Est1 allele after one generation of migration. Repeat the calculations to estimate the same allele frequency after a second generation
Answer #1

5 1 90-10 Proportion of migrants = m = = 85 Proportion of natives = * =1-m=1- 1 17 Gene frequency in mainland = Q = 0.8 = Gen

After one generation, Q = 0.8 =41 = 0.33 = 1 5+5 10 2 +90 +5 95 = 19 :: 92 = m(Q) + (1 – m)(91) 2 4 17 1 19 5+ 19 3 19 ( 5 3

**please double check the algebra; LaTex is painfully slow; thanks**